Calculate the size of a rectangular sedimentation tanks to treat a flow of 12,000 m3/d. Adopt an overflow rate of 20 m3/m2-d and detention time of 4 hours. Use two units in parallel. What will be the minimum weir length required if the weir overflow rate is not to exceed 300 m3/m-d? Also check the horizontal velocity.


1
Expert's answer
2021-09-04T01:01:15-0400

Q = 12,000 m3/d Take two units

Q = 12,000/2 = 6000 m3/d

Overflow rate (OR) = 20 m3/m2

We know

A= 6000 =300m2

Now take L/W= 2 L= 2W

Therefore 2W2=A

2W2 = 300

W = 12.24 m

L = 24.48 m

Vh= 24.48 4

= 6.12 m/hr

  = 6.12 m/hr < 9m/hr (O.K)